Sandhill Co. had the following assets on January 1, 2022. Useful Life (in years) Item Cost Purchase Date Salvage Value MachinDate Account Titles and Explanation Debit Credit Jan. 1 - Depreciation Expense Accumulated Depreciation Equipment (To record(To record sale of truck)SHOW WORK PLEASE.
Answer #1

Depreciation (year : cost - Salvage value useful life. 79000-0 machinery 10 4900/year. Forklift | 38000-O = 7600/year Truck uduaa Depreciatim machinery during - Nil Fork lift = 7600x 1 = 3800/ 6months upto June 30. Truck. = 5175. Journal Entries JanDec 31st Depreciation Accumulated $5175 depreciation - Truck. $5175 Truck was discarded on Dec 31st Book value = 44100 – 2587